# Measurement induced phase transitions in quantum raise and peel models

### v2

@@ III Numerical results
The final group, consisting of modes which apply a unitary to all of (d), (e), and (f), appear to be drawn to the frozen maximally entangled volume-law phase independent of the parameter $u$. These systems have no entanglement transition at finite $u$. For very small $u$, the equilibration time becomes very large, so we instead start the system from a state which is close to maximally entangled and observe that the system is unable to descramble the information and reach an area-law for any $u$. This maximally entangled state is reached by operating $t\gg L$ layers of random two-qubit unitary Clifford gates. We note that the modes in the first three groups are all able to descramble initially maximally entangled states when $u$ is small enough to leave the volume law.
Figure 9: Top: We show the entropy impact probability distribution for the three representative modes with phase transitions, with $u$ indicated by the color of the line. Purple corresponds to $u=0.2$ and yellow corresponds to $u=5.0$, and the color is linearly interpolated in between. In the appropriate regime, described below and in the text, the distribution function is linear on the log-log scale, indicating power-law behavior. Bottom: We show the (logarithmic) least-squares error corresponding to the best-fit to the form $P(\sigma)=A\sigma^{\tau}$. We see that the error is small in the upper phase of mode 10 and the intermediate phase of mode 20. In the area-law phase of all three modes, the entanglement is finite, so the entropy impact will be local, resulting in a poor fit to power-law behavior. Likewise, in the frozen volume-law of modes 20 and 21, the system is experiences very few measurements at the boundaries of the staircases which result in only locally affecting the entropy, again resulting in a poor fit to power-law behavior. Interestingly, for mode 21, despite transitioning directly from an area-law to a frozen volume-law, we see that for $u\lesssim u_{21}$, a power-law fit appears to have small error, indicating that the measurement impact becomes nonlocal only very near the transition point.
@@
Finally, to further connect our model to the classical RPM, we analyze a quantum entropy analogue of RPM avalanches. In particular, we compute the entropy “impact” of a measurement performed at time $t$, defined as
@@
Thus, if a measurement affects the global entropy substrate, $\sigma\rightarrow 1$, whereas if the measurement only locally affects the entropy substrate, $\sigma\rightarrow 0$. We show the results in Fig. 9. We find that the volume-law phase of mode 10 and the intermediate phase of mode 20 exhibit good fits to power-law behavior $P(\sigma)\sim\sigma^{\tau(u)}$ as characterized by the root mean square error
@@ IV Discussion and Conclusion
In this work, we consider a $(1+1)d$ quantum hybrid circuit model consisting of random Clifford unitary gates and projective measurements. For each timestep, unitary gates or projective measurements are applied to a random site depending on the local configuration of the entanglement entropy substrate according to prescribed feedback rules. This feedback mechanism, inspired by the RPM, leads to a rich phase diagram, shown in Fig. 1. We observe three distinct phases: the area-law phase, the volume-law phase with a subleading power law correction and the maximally entangled volume-law phase characterized by vanishing fluctuations. The presence of these phases are determined by the feedback mode. The phase transitions between the area-law phase and the volume-law phase with a correction term are continuous and are described universality classes distinct from those previously observed in hybrid Clifford circuits. Conversely, the phase transition to a maximally entangled frozen volume-law phase is always a first-order transition. Our findings reveal that the critical exponents associated with the former transitions closely approximate $z=1$. The frozen maximally entangled volume-law phase appears to act as a stable fixed point in the evolution for certain feedback models, resulting in a very stable configuration.
Furthermore, we compare our quantum model to the classical RPM by analyzing the distribution of avalanches, or, in the case of our model, the impact of measurements. We find that the effect of measurement on the entanglement substrate is local in the area-law phase and the maximally entangled volume-law phase, but for the volume-law phase with subleading correction term, measurements affect regions of the entropy substrate on all length scales, indicated by a power-law distribution. We find that the power-law exponent varies between -3 and -2, in agreement with the self-organized critical phase of the classical RPM. Additionally, this behavior is consistent with the volume-law phase of the hybrid random Clifford circuit.
@@ V Acknowledgements
Appendix A Results for all modes
Figure 10: We show the entanglement scaling for all modes. For each mode, we choose $L=128$ and evolve the system for $t=3000L$ timesteps followed by $2000L$ timesteps of sampling. We average over 5 repetitions of this evolution.
In this section, we show results for the entropy scaling for every mode enumerated in Fig. 3. These results are displayed in Fig. 10

### v3

@@ III Numerical results
The final model applies unitaries to all blocks except T, is drawn to the frozen maximally entangled volume-law phase independent of the parameter $u$. These systems have no entanglement transition at finite $u$. For very small $u$, the equilibration time becomes very large, so we instead start the system from a state which is close to maximally entangled and observe that the system is unable to descramble the information and reach an area-law for any $u$. This maximally entangled state is reached by operating a random $L$-qubit random Clifford gate on the state |$\ket{0}^{\otimes L}$. We note that the first three models are all able to descramble initially maximally entangled states when $u$ is small enough to leave the volume law.
@@
Finally, to further connect our model to the classical RPM, we analyze an entanglement substrate analogue of RPM avalanches. In particular, we compute the entropy “impact” of a measurement performed at time $t$, defined as
@@
Thus, if a measurement affects the global entropy substrate, $\sigma\rightarrow 1$, whereas if the measurement only locally affects the entropy substrate, $\sigma\rightarrow 0$. We show the results in Fig. 7. We find that the volume-law phase of model 1 and the intermediate phase of model 3 exhibit good fits to power-law behavior $P(\sigma)\sim\sigma^{\tau(u)}$ as characterized by the root mean square error
@@ IV Discussion and Conclusion
In this work, we consider a $(1+1)d$ quantum hybrid circuit model consisting of random Clifford unitary gates and projective measurements. For each timestep, unitary gates or projective measurements are applied to random sites depending on the local configuration of the entanglement entropy substrate according to prescribed feedback rules. This feedback mechanism, inspired by the RPM, leads to a rich phase diagram, shown in Fig. 1(c). We observe three distinct phases: the area-law phase, the volume-law phase with a subleading power law correction, and the maximally entangled volume-law phase characterized by vanishing fluctuations at late time. The presence of these phases are determined by the update rules. The phase transitions between the area-law phase and the volume-law phase with a correction term are continuous and have conformal invariance. Conversely, the phase transition to a maximally entangled frozen volume-law phase is characterized by a discontinuous jump in the entanglement entropy density in the steady states and appears to be driven by the onset of diffusive growth of the protected boundary regions. In the volume-law phase, the entanglement growth exhibits diffusive fluctuations at early times, which can be interpreted using a random walk picture for the frozen staircase region. The drift velocity decreases as the measurement rate increases and drops to zero at the transition point, leading to diffusive dynamics at the transition point.
Furthermore, we compare our quantum model to the classical RPM by analyzing the distribution of avalanches, or, in the case of our model, the impact of measurements. We find that the effect of measurement on the entanglement substrate is local in the area-law phase, but for the fluctuating volume-law phase, the measurements affect regions of the entropy substrate on all length scales, indicated by a power-law distribution. We find that the power-law exponent varies between -3 and -2, in agreement with the self-organized critical phase of the classical RPM Alcaraz et al. (2006); de Gier et al. (2004).
Interestingly, the classical RPM model can also be treated as a special non-unitary Clifford circuit. This is because there is a one-to-one correspondence between each height configuration and a non-crossing dimer configuration. In the latter, each dimer can be treated as a pair of Majorana fermions. Therefore, the RPM model can be mapped to a special Majorana fermion dynamics Nahum and Skinner (2020), which can be further described in terms of Clifford dynamics, where the measurement types and locations depend on the height profile developed in the previous time step Sang and Hsieh (2021). Currently, there is significant research on non-unitary free fermion dynamics Chen et al. (2020); Alberton et al. (2021); Turkeshi et al. (2022). Based on this perspective, we can potentially develop a deeper understanding of the phase diagram of the RPM model, most of which has been obtained numerically. We leave this for future study.